# Multi-Layer Scheduling for MoE-Based LLM Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success across a wide range of tasks, but serving them efficiently at scale remains a critical challenge due to their substantial computational and latency demands. While most existing inference frameworks rely on simple scheduling strategies such as First-Come-First-Serve (FCFS) at the engine level and Round-Robin (RR) at the scheduler or coordinator level, they often fail to fully utilize system resources and may suffer from issues such as head-of-line blocking and load imbalance. Recent advances in Mixture-of-Experts (MoE) models have also introduced new challenges in scheduling arising from expert parallelism and routing complexity. This research proposes a multi-layer scheduling framework tailored for MoE-based LLM serving. It targets scheduling at three levels: request-level, engine-level, and expert-level. At the request level, we explore algorithms such as Shortest-Job-First (SJF) and priority-aware aging to improve throughput and reduce latency. At the engine level, we design load-aware dispatching strategies that account for the current prefix token load, KV cache utilization, and user stickiness to achieve better resource matching. At the expert level, we focus on alleviating expert hotspots and strategically placing inter-layer expert dependencies to balance load and improve routing efficiency. Extensive experimental results from more than 100 experiments conducted under diverse workload distributions show that our approach consistently outperforms the state-of-the-art inference framework vLLM, achieving up to 17.8% reduction in Time To First Token (TTFT) latency and 13.3% reduction in Time-Per-Output-Token (TPOT) latency.
@@ II-C Motivation
Currently, mainstream LLM serving systems such as vLLM [12] and SGLang [32] mainly adopt service frameworks similar to Figure 1 designed for traditional dense models, lacking specific optimizations for the characteristics of MoE models. In dense model scenarios, scheduling is relatively simple, often performed on a single machine or single GPU using FCFS, or via simple Round-Robin (RR) strategies across multiple inference engines in multi-GPU settings. While these approaches are simple and easy to implement, they are insufficient for MoE serving scenarios: First, engine-level RR scheduling is unaware of each engine’s current load status, such as KV Cache usage, which can result in some engines becoming overloaded while others remain underutilized, leading to suboptimal throughput and latency.
@@ III-D1 Problem Formulation
be the activation matrix, where $A_{i,j}$ denotes the activation intensity of expert $j$ at layer $i$. Let $g<m$ be the number of available GPUs, and assume each GPU hosts approximately $m/g$ experts.
@@ III-D2 MILP Formulation
The parameter $D$ specifies the maximum deviation allowed between the ideal balanced activation load $L_{i}$ and the actual load $L_{i,p}$ assigned to each GPU at layer $i$. A smaller value of $D$ enforces strict load balance, ensuring that every GPU receives nearly equal activation volume at each layer. However, such strictness may require excessive expert relocation and can interfere with desirable co-location of affinity-linked experts. Conversely, a larger $D$ relaxes the balance constraint, allowing certain GPUs to host heavier activation loads in exchange for improved communication locality.
@@
Define same-part indicator $s_{j,k,p}\in\{0,1\}$ to denote whether experts $j$ and $k$ are both placed on GPU $p$. This is linearized using:
@@ III-D3 Implementation
Inputs: affinity matrix $\mathcal{M}$,  activation matrix $A$,  GPU index $k$,  number of GPUs $g$,  step interval $\tau$.
@@ VI Related Work
Overall, while these approaches demonstrate the importance of moving beyond traditional FCFS and RR baselines, most optimizations focus on a single layer—either request scheduling, engine dispatching, or expert routing. In contrast, our work explores scheduling optimizations at multiple layers of LLM serving, including the request, engine, and expert levels. Specifically, Gimbal combines request-level SJF with, engine-level load-aware scheduling (considering prefix tokens, KV Cache usage, and user stickiness), and expert-level hotspot mitigation with dependency-aware placement, aiming to holistically improve MoE-based LLM serving efficiency.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success across a wide range of tasks, but serving them efficiently at scale remains a critical challenge due to their substantial computational and latency demands. While most existing inference frameworks rely on simple scheduling strategies such as First-Come-First-Serve (FCFS) at the engine level and Round-Robin ($R$) at the scheduler or coordinator level, they often fail to fully utilize system resources and may suffer from issues such as head-of-line blocking and load imbalance. Recent advances in Mixture-of-Experts (MoE) models have also introduced new challenges in scheduling arising from expert parallelism and routing complexity. This research proposes a multi-layer scheduling framework tailored for MoE-based LLM serving. It targets scheduling at three levels: request-level, engine-level, and expert-level. At the request level, we explore algorithms such as Shortest-Job-First (SJF) and priority-aware aging to improve throughput and reduce latency. At the engine level, we design load-aware dispatching strategies that account for the current prefix token load, KV cache utilization, and user stickiness to achieve better resource matching. At the expert level, we focus on alleviating expert hotspots and strategically placing inter-layer expert dependencies to balance load and improve routing efficiency. Extensive experimental results from more than 100 experiments conducted under diverse workload distributions show that our approach consistently outperforms the state-of-the-art inference framework vLLM, achieving up to 17.8% reduction in Time To First Token (TTFT) latency and 13.3% reduction in Time-Per-Output-Token (TPOT) latency.
@@ II-C Motivation
Currently, mainstream LLM serving systems such as vLLM [12] and SGLang [32] mainly adopt service frameworks similar to Figure 1 designed for traditional dense models, lacking specific optimizations for the characteristics of MoE models. In dense model scenarios, scheduling is relatively simple, often performed on a single machine or single GPU using FCFS, or via simple Round-Robin ($R$) strategies across multiple inference engines in multi-GPU settings. While these approaches are simple and easy to implement, they are insufficient for MoE serving scenarios: First, engine-level $R$ scheduling is unaware of each engine’s current load status, such as KV Cache usage, which can result in some engines becoming overloaded while others remain underutilized, leading to suboptimal throughput and latency.
@@ III-D1 Problem Formulation
be the activation matrix, where $A_{i,j}$ denotes the activation intensity of expert $j$ at layer ii. Let $g<m$ be the number of available GPUs, and assume each GPU hosts approximately $m/g$ experts.
@@ III-D2 MILP Formulation
The parameter $D$ specifies the maximum deviation allowed between the ideal balanced activation load $L_{i}$ and the actual load $L_{i,p}$ assigned to each GPU at layer ii. A smaller value of $D$ enforces strict load balance, ensuring that every GPU receives nearly equal activation volume at each layer. However, such strictness may require excessive expert relocation and can interfere with desirable co-location of affinity-linked experts. Conversely, a larger $D$ relaxes the balance constraint, allowing certain GPUs to host heavier activation loads in exchange for improved communication locality.
@@
Define same-part indicator $s_{j,k,p}\in\{0,1\}$ to denote whether experts $j$ and $k$ are both placed on GPU pp. This is linearized using:
@@ III-D3 Implementation
Inputs: affinity matrix $\mathcal{M}$,  activation matrix AA,  GPU index $k$,  number of GPUs $g$,  step interval $\tau$.
@@ VI Related Work
Overall, while these approaches demonstrate the importance of moving beyond traditional FCFS and $R$ baselines, most optimizations focus on a single layer—either request scheduling, engine dispatching, or expert routing. In contrast, our work explores scheduling optimizations at multiple layers of LLM serving, including the request, engine, and expert levels. Specifically, Gimbal combines request-level SJF with, engine-level load-aware scheduling (considering prefix tokens, KV Cache usage, and user stickiness), and expert-level hotspot mitigation with dependency-aware placement, aiming to holistically improve MoE-based LLM serving efficiency.